# Un-Attributability: Computing Novelty from Retrieval & Semantic Similarity

### v1

@@ Abstract
Understanding how language‑model outputs relate to the pretraining corpus is central to studying model behavior. Most training‑data attribution (TDA) methods ask which training examples causally influence a given output, often using leave‑one‑out tests. We invert the question: which outputs cannot be attributed to any pretraining example? We introduce un-attributability as an operational measure of semantic novelty: an output is novel if the pretraining corpus contains no semantically similar context. We approximate this with a simple two-stage retrieval pipeline: index the corpus with lightweight GIST embeddings, retrieve the top‑$n$ candidates, then rerank with ColBERTv2. If the nearest corpus item is less attributable than a human-generated text reference, we consider the output of the model as novel. We evaluate on SmolLM and SmolLM2 and report three findings: (1) models draw on pretraining data across much longer spans than previously reported; (2) some domains systematically promote or suppress novelty; and (3) instruction tuning not only alters style but also increases novelty. Reframing novelty assessment around un-attributability enables efficient analysis at pretraining scale. We release $\sim$20 TB of corpus chunks and index artifacts to support replication and large‑scale extension of our analysis at https://huggingface.co/datasets/stai-tuebingen/faiss-smollm.
@@ 1 Introduction
Large language models (LLMs) now power chatbots, copilots, and autonomous agents. Understanding how language model outputs relate to their pretraining corpora is central to studying model behavior and generalization. A key question is whether an output is novel – not traceable to memorizing training data. Measuring novelty reveals when models generalize beyond what they have seen, signals compositional generalization, assessing the true zero-shot behavior – informing debates about provenance and intellectual property. Thus, computing novelty is both a technical problem and a prerequisite for interpreting what LLMs actually learn.
Training data attribution (TDA) methods address related questions by tracing model behavior back to specific data (Hammoudeh & Lowd, 2024; Deng et al., 2025). Two approaches dominate (Chang et al., 2025): causal influence methods measure a training sample’s leave-one-out effect (Koh & Liang, 2017) but do not scale to trillion-token corpora; factual attribution methods scale but rely on lexical matches between outputs and training text (Liu et al., 2025b; a; Wang et al., 2025c), not being robust to many simple variations. These approaches have taught us much about what can be definitively attributed. Yet many consequential questions hinge on the opposite: what cannot be attributed.
We address this gap by inverting the usual TDA question and focusing on un-attributability. Rather than asking which training samples caused an output, we ask which outputs cannot be attributed to the pretraining corpus. We treat such outputs as novel. Our test is strict: the pretraining data must contain neither lexical overlaps nor semantically similar contexts with the generation (Figure 1).
We operationalize this test using best-match semantic similarity between a generation and the full pretraining corpus. Concretely, we compute GIST embeddings (Solatorio, 2024) over corpus chunks, build a vector index, retrieve the top-$n$ candidates, and re‑rank them with ColBERTv2 (Santhanam et al., 2022). To calibrate attribution scores, we also evaluate held-out human-written references that are guaranteed not to appear in the pretraining data. If even a strong ranker such as ColBERTv2 assigns lower similarity to an LLM output than to such known-novel human text, we deem it as novel. This procedure offers greater coverage and auditability than lexical matching, is model‑ and task‑agnostic, and is lightweight enough to run on full pretraining corpora. It thus offers a scalable way to collect evidence of what is not attributable.
@@
We apply the test to SmolLM (Allal et al., 2024) and SmolLM2 (Allal et al., 2025), two LLMs with open pretraining corpora. Our analysis reveals surprising patterns missed by previous lexical methods (McCoy et al., 2023; Merrill et al., 2024). First, both models draw on pretraining data over much longer sequences than previously reported. Second, novelty varies systematically by task domain. Third, embedding-based novelty estimates are stable under style shifts from instruction tuning; after accounting for these shifts, instruction tuning substantially increases novelty. These results suggest that instruction tuning shapes not only style but also compositional generation behavior.
Our contributions are conceptual and practical. Conceptually, we invert the goal of TDA from what is attributable to what is definitively not attributable. Practically, we present a lightweight test for model novelty based on un-attributability and conduct a large-scale study on SmolLM and SmolLM2. We also release the corresponding indices and corpus chunks of SmolLM and SmolLM2 to support replication and extension.
Figure 1: Embedding similarity is more robust to long or paraphrased texts than N-gram similarity. Comparison of similarity measured with N-gram overlap (left) and embedding cosine similarity (right) with increasing sequence length. The similarity is measured between a model generation and its top 2 closest semantic matches in the pretraining corpus retrieved using our test. Both training excerpts convey the same information as the generation, but lexical overlap fails to recognize this with larger N-grams, whereas embeddings remain robust.
@@ 2 Related Work
We situate our work in relation to fields studying the capabilities of a model in relation to the training data, namely, training data attribution (TDA), causal and factual, as well as memorization research.
Causal TDA.
Causal TDA attributes model behavior to a training sample by the treatment effect that including the sample in the training dataset has on the observed behavior, also known as the leave-one-out (LOO) effect. Since LOO is infeasible to compute explicitly across datasets, the field of TDA studies approximations of LOO, where influence functions (Hampel, 1974) (IF) are most prominent. Method work studies how to adapt IFs to deep models, where convexity assumptions break, computational and memory costs explode, resulting in a variety of IF adaptations (Koh & Liang, 2017; Schioppa et al., 2022; Chang et al., 2025; Grosse et al., 2023; Park et al., 2023; Pruthi et al., 2020; Wang et al., 2025a; Choe et al., 2024; Xia et al., 2024). However, IFs are known to be fragile when assumptions are not met (Bae et al., 2022; Basu et al., 2021; Epifano et al., 2023; Nguyen et al., 2023). Another line of work estimates LOO through unrolled differentiation through the training process, bypassing model convergence assumptions and considering different optimizer algorithms (Hara et al., 2019; Bae et al., 2024; Wang et al., 2025b; Ilyas & Engstrom, 2025). While both approaches approximate LOO in deep models like LLMs, their memory and computational demands limit scalability to large pretraining corpora, requiring workarounds. For instance, Guo et al. (2021) reduce costs by retrieving candidate samples via $k$-nearest neighbors in embedding space, while Grosse et al. (2023) rely on TF-IDF filtering to preselect potentially influential samples. These approaches assume that influential samples must lie close in some representation space, but this is not guaranteed (Hu et al., 2024) when not explicitly trained for it, like in Yeh et al. (2018); Sun et al. (2025). We invert this logic: when the nearest samples in representation space are dissimilar from the generation, we interpret this un-attributability as novelty – evidence that the model’s output is composed rather than reused.
Factual TDA.
Factual TDA attributes LLM outputs back to the training data to identify what grounds the output (Chang et al., 2025). For this task, retrieval methods relying on lexical overlap like BM25 and $n$-gram overlap (Liu et al., 2025a; Merrill et al., 2024; McCoy et al., 2023; Liu et al., 2025b; Gottesman et al., 2025; Peng et al., 2023; Wang et al., 2025c) offer a strong baseline and are often better suited than causal TDA (Akyürek et al., 2022; Chang et al., 2025). Additionally, lexical overlap-based retrieval methods are considerably more lightweight than causal TDA, allowing them to scale efficiently to large datasets such as pretraining corpora. However, lexical overlap is an overly strict criterion, as it underestimates the possibility that a model learns facts from paraphrased documents. For factual attribution, or in our case, evaluating novelty by un-attributability, it is more appropriate to consider a semantic representation of the generation and measure semantic similarity.
@@
Closely related are studies of memorization and membership inference. Memorization work (Wu et al., 2025; Feldman & Zhang, 2020) typically investigates whether specific samples can be elicited verbatim from a model, whereas our notion of novelty captures whether the underlying information is present in the corpus, even if paraphrased. Membership inference attacks (MIA) (Puerto et al., 2025; Mesana et al., 2025; Zhang et al., 2025; 2024) instead ask whether a particular example was part of pretraining, often in adversarial settings. While informative for data privacy, MIAs do not address the broader question of how models generate text not attributable to their training data. Our novelty measure, therefore, complements both attribution and memorization/MIA, providing a new perspective on generalization.
3 Defining novelty as what is Un-attributable
Figure 2: Pipeline for scoring the novelty of an LLM output $q$. We test whether $q$ is unattributable to the pretraining corpus – our operational definition of novelty. Stage 0 (one-time): Chunk the corpus, compute L2-normalized GIST (Solatorio, 2024) embeddings, and build a cosine-similarity FAISS (Douze et al., 2024) index. Stage 1: Embed $q$ with GIST and retrieve the top-$n$ nearest corpus chunks. Stage 2: Rerank retrieved candidates with ColBERTv2 (Santhanam et al., 2022) at multiple chunk sizes. The novelty score is the median, over $q$’s chunks, of the ColBERTv2 similarity to the best retrieved chunk, normalized by the sequence length and corresponding baseline score.
We propose a retrieval-based test for semantic novelty in LLM outputs, analyzing generalization with respect to the training data. The central challenge is the scale of pretraining corpora. To make the problem tractable, we focus on un-attribution: showing that no close semantic match to the output exists in the corpus. This reduces the task to large-scale retrieval. If no close match is found, we deem the output novel rather than attributable to the corpus. We use a two-stage pipeline (Figure 2). Stage 1 performs initial retrieval over GIST embeddings with FAISS (Douze et al., 2024); Stage 2 reranks candidates with ColBERTv2 (Santhanam et al., 2022). Algorithm 1 details the procedure, and we will release our code upon publication.
3.1 Stage 1: building the vector index and initial retrieval
We first build a cosine-similarity-based FAISS index (Douze et al., 2024) $I_{\mathrm{FAISS}}$ to enable efficient search over the entire pretraining corpus. Because FAISS operates on vectors, we embed corpus chunks with the GIST model (Solatorio, 2024). GIST embeddings are compact, reducing storage costs; yet expressive, as evidenced by strong MTEB performance (Muennighoff et al., 2023; Enevoldsen et al., 2025), allowing accurate attribution at pretraining scale. We chunk text at the GIST model’s maximum length (512 tokens) and L2-normalize all embeddings so inner products equal cosine similarity. [^1] We publish the searchable index and corresponding corpus chunks at https://huggingface.co/datasets/stai-tuebingen/faiss-smollm. At query time, we embed and L2-normalize the LLM output $q$, then retrieve the top-$n=100$ nearest corpus chunks, balancing recall and compute. Please refer to Appendix A for additional analysis about this choice.
3.2 Stage 2: Re-ranking using ColBERTv2 for fine-grained analysis
We then rerank the top 100 candidates with finer-grained similarity via ColBERTv2 (Santhanam et al., 2022). We chunk both the LLM output and the retrieved candidates at one or more sizes and rerank with ColBERTv2, which operates on context-dependent token embeddings and is widely used for reranking (Zhao et al., 2024). Varying chunk size lets us examine how length affects novelty.
While ColBERTv2 scores are accurate, they have two drawbacks: (1) without a reference, they are hard to interpret; and (2) because they sum over query tokens, they are length-biased. We address both by (a) normalizing scores by query-token count and (b) comparing each output to a human-written domain-matched baseline $b$, if available. The resulting similarity ratio is interpretable: values <1<1 indicate $q$ is more novel than the baseline; values >1>1 indicate greater attributable overlap. This mechanism allows us to compare model novelty relative to human-written text.
@@ 3.2 Stage 2: Re-ranking using ColBERTv2 for fine-grained analysis
Algorithm 1 Novelty test with retrieval, reranking, and baseline-normalized scoring
$I_{\text{FAISS}}$ (FAISS index over L2-normalized GIST embeddings), $q$ (LLM output), $b$ (baseline text), $\phi_{\mathrm{GIST}}$ (GIST embedder), $\mathrm{ColBERT}$ (ColBERTv2 reranker), $n>0$ (initial retrieval size), $k>0$ (chunk size in tokens)
Stage 1: Initial retrieval (cosine similarity via FAISS)
$D_{\!q}\leftarrow\text{kNN}(I_{\text{FAISS}},\phi_{\mathrm{GIST}}(q),n)$
$D_{\!b}\leftarrow\text{kNN}(I_{\text{FAISS}},\phi_{\mathrm{GIST}}(b),n)$
Stage 2: Reranking and normalized scoring
$Q^{(k)}\leftarrow\text{chunk}(q,k)$; $B^{(k)}\leftarrow\text{chunk}(b,k)$
$C^{(k)}_{\!q}\leftarrow\bigcup_{d\in D_{\!q}}\text{chunk}(d,k)$; $C^{(k)}_{\!b}\leftarrow\bigcup_{d\in D_{\!b}}\text{chunk}(d,k)$
Define length-normalized best-match score for chunk $x$:
$\qquad\widetilde{s}(x,C)\;=\;\displaystyle\frac{\max_{c\in C}\ \mathrm{ColBERT}(x,c)}{|x|}$ ($|x|$ = # query tokens used by ColBERT)
Compute baseline normalizer (robust central tendency):
$\mu_{B}^{(k)}\leftarrow\operatorname{mean}\big(\{\,\widetilde{s}(b^{(k)},C^{(k)}_{\!b})\;:\;b^{(k)}\in B^{(k)}\,\}\big)$
Compute per-chunk similarity ratios for $q$:
$R^{(k)}\leftarrow\big[\,\widetilde{s}(q^{(k)},C^{(k)}_{\!q})\,/\,\mu_{B}^{(k)}\ :\ q^{(k)}\in Q^{(k)}\,\big]$
Output: summary novelty score and diagnostics
$N^{(k)}\leftarrow\operatorname{median}(R^{(k)})$ $\triangleright$ Novelty/attributability ratio for $k$; <1<1 = more novel than baseline
return $N^{(k)},\,R^{(k)}$

### v2

# LLM generation novelty through the lens of semantic similarity

@@ Abstract
Generation novelty is a key indicator of an LLM’s ability to generalize, yet measuring it against full pretraining corpora is computationally challenging. Existing evaluations often rely on lexical overlap, failing to detect paraphrased text, or do not consider the full pretraining corpus. We frame novelty as a semantic retrieval problem. This framing enables us to address novelty with modern embedding and indexing pipelines, allowing for efficient analysis at pre-training scale. Specifically, we propose a three-stage framework that retrieves semantically similar samples, reranks them at varying subsequence lengths, and calibrates scores using a human novelty reference for interpretability. We apply this framework to the SmolLM model family and report three key findings: (1) models draw on pretraining data across much longer sequences than previously reported; (2) some task domains systematically promote or suppress generation novelty; and (3) instruction tuning not only alters style but also increases novelty. These results highlight the value of semantic novelty analysis for studying generalization. To support reproducibility and further research, we release $\sim$20 TB of corpus chunks and index artifacts at https://huggingface.co/datasets/stai-tuebingen/faiss-smollm.
LLM generation novelty through the lens of semantic similarity
  Philipp Davydov   Ameya Prabhu   Matthias Bethge  Elisa Nguyenthanks: Equal Supervision   Seong Joon Oh11footnotemark: 1 Tübingen AI Center, University of Tübingen
@@ 1 Introduction
Figure 1: Efficient retrieval allows for pretraining-scale semantic novelty analysis. Given a generation, we employ retrieval and reranking pipelines to identify semantically similar samples in the pretraining corpus to analyze generation novelty. Higher similarity indicates lower novelty. While the first-ranked sample is a lexical match, the second-ranked is a semantic match. As text length increases, $n$-gram overlap drops to zero, falsely suggesting high novelty for paraphrased content. In contrast, embedding similarity correctly identifies the shared meaning, providing a more reliable measure of novelty.
@@
Large language models (LLMs) now power chatbots, copilots, and autonomous agents with applications across various domains. Their adoption hinges on an implicit assumption: LLMs are capable of generalizing beyond their training data to generate relevant and novel outputs in response to user prompts. Generation novelty provides a useful signal of this capability, indicating whether model outputs extend beyond patterns observed during training. Novelty signals compositional generalization and allows for the assessment of true zero-shot behavior – informing debates about provenance and intellectual property. Thus, measuring a model’s generation novelty is a prerequisite for interpreting what LLMs actually learn.
However, analyzing the novelty of LLM generations is a non-trivial task. There is no single, clearly defined notion of novelty, which opens up the question: Is a generation already novel if it does not appear verbatim in the pretraining data? More broadly, what does it mean for a generation to be “reused” from the training data? At the same time, the sheer scale of modern pretraining corpora makes comparison computationally challenging, so that novelty analysis is both a conceptual and a technical problem.
Existing approaches to studying LLM generation novelty mainly address these questions in two directions. One line of work relies on efficient textual overlap metrics (McCoy et al., 2023; Merrill et al., 2024; Padmakumar et al., 2025), which scale well but naturally focus on verbatim reuse and fail to account for paraphrases or stylistic variation. A second line of work investigates novelty within specific domains, such as scientific ideas or biomedical publications (Peng et al., 2025; Ai et al., 2025; Wang et al., 2025c), typically using semantic similarity measured against a curated reference corpus. While effective, these approaches do not account for reuse from other parts of the broader pretraining corpus. Together, these limitations highlight the need for novelty measures that go beyond verbatim overlap at the scale of LLM pretraining corpora.
We address this need by proposing a three-stage framework for analyzing an LLM’s generation novelty at scale through the lens of semantic similarity. We argue that a generation should be called novel only if it does not semantically reproduce the training data (cf. Figure 1). From this perspective, “reused” information is not limited to verbatim overlap, but includes paraphrases and reformatting.
Our framework combines semantic retrieval (stage 1) and reranking (stage 2) (Santhanam et al., 2022; Li et al., 2025) to compare LLM generations against their pretraining corpora. However, similarity scores and their relative differences are not directly interpretable as measures of novelty. They are affected by biases in the retrieval pipeline, such as a preference for shorter documents and information located early within documents (Fayyaz et al., 2025; Zhou et al., 2025). Because of this, identical scores can reflect different degrees of novelty, especially when comparing different text lengths or domains. Hence, the third stage of our framework calibrates the raw semantic similarity scores using a baseline of held-out, human-written reference text. By calibrating the scores, we mitigate artifacts of the retrieval pipeline and enable meaningful comparison of scores. Finally, we aggregate these scores into a novelty profile which characterizes the generation novelty of a model for a specific dataset and generation text length. This procedure is model- and task-agnostic, and lightweight enough to run on full pretraining corpora. It thus offers a scalable way to assess an LLM’s generation novelty.
@@
We demonstrate the utility of our framework through empirical analyses on SmolLM (Allal et al., 2024) and SmolLM2 (Allal et al., 2025), two LLMs with open pretraining corpora. Our analysis reveals unexpected patterns missed by previous lexical methods (McCoy et al., 2023; Merrill et al., 2024). First, both models draw on pretraining data over much longer sequences than previously reported (Merrill et al., 2024). Second, novelty varies systematically by task domain. Third, embedding-based novelty estimates are stable under style shifts from instruction tuning; after accounting for these shifts, instruction tuning substantially increases novelty. These results suggest that instruction tuning shapes not only style but also compositional generation behavior.
Our contributions are as follows:
  • We present a semantic similarity–based framework for measuring LLM generation novelty at scale, combining retrieval, reranking, and baseline-calibration to enable comparison against full pretraining corpora while reducing sensitivity to stylistic variation.
  • We empirically analyze generation novelty in SmolLM and SmolLM2, uncovering long-range reuse patterns, task-dependent variation, and the impact of instruction tuning beyond stylistic effects.
  • We also release the corresponding indices and corpus chunks of SmolLM and SmolLM2 to support replication and extension.
@@ 2 Related Work
Novelty measurement in LLMs.
Prior work on LLM generation novelty has predominantly relied on textual overlap–based measures, particularly $n$-gram comparisons. McCoy et al. (2023) introduce $n$-novelty, defining novelty as the absence of copied text from the training data and quantifying the proportion of non-overlapping $n$-grams in model outputs. Building on this, Merrill et al. (2024) additionally analyze the probability of generating training $n$-grams, while Padmakumar et al. (2025) propose novelty as the harmonic mean of $n$-novelty and generation quality assessed by LLM judges. Wang et al. (2025b) further extend this line of work by introducing task-grams, which capture task-specific $n$-gram co-occurrences in output and corpus. The task relation introduces a semantic dimension to the $n$-gram-based analysis. Despite these extensions, all of these approaches rely on surface-level textual overlap and may treat paraphrases or stylistic variation as novel. We instead define novelty via semantic similarity, enabling analysis beyond surface-form variation. Prior work that emphasizes semantic novelty typically focuses on specific domains, such as scientific ideas or biomedical text (Ai et al., 2025; Peng et al., 2025; Wang et al., 2025c). These approaches also rely on embedding-based similarity, but assess novelty relative to selected reference documents tailored to the target domain. In contrast, our work studies generation novelty with respect to the entire pretraining corpus, independent of the prompt.
@@
Memorization work (Wu et al., 2025; Feldman and Zhang, 2020) investigates whether specific samples can be elicited verbatim from a model to understand if the model has memorized them. In contrast, our notion of novelty captures whether the underlying information is present in the corpus, even if paraphrased. Membership inference attacks (MIA) (Puerto et al., 2025; Mesana et al., 2025; Zhang et al., 2025, 2024) instead ask whether a particular example was part of pretraining, often in adversarial settings. While informative for questions like data privacy, MIAs do not address the broader question of how models generate text that are not part of their training data. Our novelty measure, therefore, complements both attribution and memorization/MIA, providing a new perspective on generalization.
Further discussion of prior work relating model generations to pretraining data is provided in Appendix A.
3 Novelty analysis framework
This section presents our framework for studying LLM generation novelty. We introduce the conceptual definition of semantic novelty and propose a framework that measures calibrated semantic novelty at LLM pretraining scale.
3.1 Conceptual Definition of Semantic Novelty
We define generation novelty as the degree to which an LLM’s output expresses information or patterns not readily present in its pretraining corpus. Unlike lexical novelty, which is typically measured via $n$-gram non-overlap (McCoy et al., 2023; Merrill et al., 2024), our definition focuses on semantic novelty:
$\min_{d\in\mathcal{C}}\cos(\phi(y),\phi(d)),y\sim f_{\theta}(x)$ (1)
where $\mathcal{C}$ is the training corpus, $y$ is the LLM $f_{\theta}$’s output to a prompt $x$, and $\phi$ is an embedding function. In other words, semantic novelty is defined as the minimum cosine similarity between a semantic representation of the LLM generation and documents of the pretraining corpus. Under this definition, an output’s novelty is determined by its content, rather than its surface form. Hence, this definition serves to distinguish between reproduction (generating sequences that exist semantically in the corpus) and composition (generating sequences that are internally coherent and correct, yet semantically distinct from any specific training document).
By aggregating semantic signals across a dataset, we can characterize the novelty profile of a specific model configuration or training regime. Crucially, we treat novelty not as a binary state, but as a continuous spectrum. We do not claim that a “novel” generation has no relationship to the training data; rather, we measure the extent to which the generated content deviates from the most similar semantic neighbors available in the corpus.
3.2 A Framework for Analysing Semantic Novelty
@@
Algorithm 1 Novelty Framework with retrieval, reranking and baseline-normalized scoring
$\mathcal{C}$ (Corpus), $\mathcal{Q}$ (Model generations), $\mathcal{H}$ (Human references), $\mathcal{R}$ (Ranker), $\mathcal{S}$ (Re-ranker), $K$ (Set of chunk sizes $\{k_{1},k_{2},\dots\}$)
// Stage 1: Identification of Local Candidate Pools
for each document $d\in\mathcal{Q}\cup\mathcal{H}$ do
  $P_{d}\leftarrow\mathcal{R}(d,\mathcal{C},n)$ $\triangleright$ Retrieve $n$ passages once per document using Eq. 1
end for
// Stage 2: Multi-Scale Evaluation and Calibration
for each chunk size $k\in K$ do
  $\text{Scores}_{H}\leftarrow\emptyset,\text{Scores}_{Q}\leftarrow\emptyset$
// Determine the "Semantic Noise Floor" for this configuration
  for each document $h\in\mathcal{H}$ do
   $Chunks_{h}\leftarrow\text{chunk}(h,k)$
   $S_{h}\leftarrow\{\max_{p\in P_{h}}\mathcal{S}(c,p)\mid c\in Chunks_{h}\}$
   $\text{Scores}_{H}\leftarrow\text{Scores}_{H}\cup S_{h}$
  end for
  $\mu_{H}^{(k)}\leftarrow\text{mean}(\text{Scores}_{H})$ $\triangleright$ Stable calibrator for domain and length $k$
// Stage 3: Calculate Calibrated Similarity Scores for the Model
  for each document $q\in\mathcal{Q}$ do
   $Chunks_{q}\leftarrow\text{chunk}(q,k)$
   $S_{q}\leftarrow\{\max_{p\in P_{q}}\mathcal{S}(c,p)\mid c\in Chunks_{q}\}$
   $R_{q}\leftarrow\{s/\mu_{H}^{(k)}\mid s\in S_{q}\}$ $\triangleright$ Normalize by human distribution
   $\text{Scores}_{Q}\leftarrow\text{Scores}_{Q}\cup R_{q}$
  end for
  $N^{(k)}\leftarrow\text{median}(\text{Scores}_{Q})$ $\triangleright$ Final novelty profile for chunk size $k$
end for
return $\{(k,N^{(k)})\mid k\in K\}$ $\triangleright$ The novelty profile of the model
Our framework operationalizes the measurement of semantic novelty through a model-agnostic, three-stage paradigm applied at the dataset scale. This approach is inspired by standard retrieval-augmented generation (RAG) and information retrieval architectures, where a two-stage process of ranking and re-ranking is the established procedure for balancing search efficiency with semantic precision (Li et al., 2025). Since our framework addresses generation novelty, we further append a third stage that mitigates potential skewness artifacts of the retrieval pipeline, arising from sequence length or domain-specific density of the LLM generations. This stage calibrates the novelty scores with respect to a matched human-level novelty reference so that scores are comparable and interpretable.
The framework aims to characterize the novelty profile $\mathcal{N}_{\mathcal{Q}}(k)$ of a model distribution $\mathcal{Q}$ relative to a domain-matched human distribution $\mathcal{H}$, for each chunk size $k$. By evaluating across multiple chunk sizes $k$, we can observe how specific model configurations (e.g., scale or alignment) influence novelty. The framework is summarized in Algorithm 1. We describe each stage in the following:
  1. 1.
    Local Candidate Pool Identification. For every generation in the model distribution $q\in\mathcal{Q}$ and corresponding baseline from the human distribution $h\in\mathcal{H}$, first identify a local candidate pool $D$ within the pretraining corpus $\mathcal{C}$. To this end, a coarse-grained ranker $\mathcal{R}$ retrieves the top-$n$ passages that exhibit the highest potential semantic overlap with the full document. This stage acts as a high-recall filter, ensuring that any potential semantic neighbors are captured once per document, providing a computationally efficient search space for subsequent multi-scale analysis.
  2. 2.
    Multi-Scale Semantic Re-ranking. To distinguish between short-range lexical reuse and long-range compositional novelty, evaluate text at multiple chunk sizes $k$. For a given $k$, decompose the generation and its corresponding candidate pool into smaller fragments. A high-precision re-ranker $\mathcal{S}$ then computes the maximum similarity between each chunk and its respective pool. This yields a set of raw similarity scores for the entire distribution. This two-stage process ensures that the scoring is fine-grained and robust to paraphrasing while remaining tractable at the scale of trillion-token corpora.
  3. 3.
    Calibrated Distributional Normalization. As human and model outputs differ in length and structure, a 1:1 element-wise comparison between individual chunks is often impossible: Given one human baseline text for each model generation, they ultimately are split into a different number of chunks for small enough $k$. Instead, we use the human distribution $\mathcal{H}$ to establish a stable, domain-specific calibration constant $\mu_{H}^{(k)}$ for each chunk size. This constant represents the semantic noise floor, i.e., the level of similarity expected from novel, held-out human text within that domain, configuration, and chunk size. Further motivation and details on the baseline calibration are provided in Appendix C.
The final calibrated similarity score for the model distribution is calculated by normalizing the model’s raw scores by the semantic noise floor constant. By aggregating these ratios, we arrive at a novelty profile that allows for rigorous comparison across different model architectures or training regimes. This normalization step ensures that any observed trends are not artifacts of the retrieval pipeline or domain-specific redundancies, but true reflections of the model’s divergence from natural human patterns of information reuse.